# TropNNC: Structured Neural Network Compression Using Tropical Geometry

### v1

@@ A.6 Proof of Proposition 14
Consider any vertex $\mathbf{u}$ of $P_{j}$. This vertex can be written as the sum of generators $c_{ji}(\mathbf{a}_{i}^{T},b_{i})$, for some subset $I_{j+}^{\prime}\subset I_{j+}$. Thus,
$u=\sum_{i\in I_{j+}^{\prime}}c_{ji}(\mathbf{a}_{i}^{T},b_{i})$
For every generator $c_{ji}(\mathbf{a}_{i}^{T},b_{i})$ that belongs to cluster k, define
$x_{ji}=arg\min_{x}\left\{c_{ji}(\mathbf{a}_{i}^{T},b_{i})-x\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\}$
@@
$\sum_{j=1}^{m}\left[|\tilde{C}_{jk}|(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})-\sum_{i\in I_{jk}}|C_{ji}|(\mathbf{a}_{i}^{T},b_{i})\right]^{2}.$
@@
By hypothesis, for every cluster $k$, the vectors of the set $\{(\mathbf{a}_{i}^{T},b_{i})|i\in I_{k}\}$ form pair-wise acute angles. Every vector of the set of vectors $\{\sum_{i\in I_{jk}}|c_{ji}|(\mathbf{a}_{i}^{T},b_{i})|j\in[m]\}$ lies inside the cone of the set of vectors $\{(\mathbf{a}_{i}^{T},b_{i})|i\in I_{k}\}$. It is easy to verify that the representative $|\tilde{c}_{jk}|(\tilde{\mathbf{a}}_{k},\tilde{b}_{k})$ that the iterative algorithm produces lies inside the cone of the set of vectors $\{\sum_{i\in I_{jk}}|c_{ji}|(\mathbf{a}_{i}^{T},b_{i})|j\in[m]\}$, which is a subset of the cone of the set of vectors $\{(\mathbf{a}_{i}^{T},b_{i})|i\in I_{k}\}$. Thus, the representative forms an acute angle with every vector of the set $\{(\mathbf{a}_{i}^{T},b_{i})|i\in I_{k}\}$, and thus $x_{ji}\geq 0$
Define $I_{jk}^{\prime}=I_{k}\cap I_{j+}^{\prime}$. For every cluster k, define
$\tilde{x}_{jk}=\sum_{i\in I_{jk}^{\prime}}x_{ji}$
@@
Thus, for every cluster $k$, the point $\tilde{x}_{jk}\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ lies inside the segment $[\text{O},\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$], and thus belongs to $\tilde{P_{j}}$.
For every vertex $u$, we choose to compare it with the vertex
$\sum_{k\in C_{j+}}\tilde{x}_{jk}\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\in\tilde{P_{j}}$
@@
$\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})=\sum_{i\in I_{jk}}c_{ji}(\mathbf{a}_{i}^{T},b_{i})+l_{jk}=|I_{jk}|\frac{\sum_{i\in I_{jk}}c_{ji}(\mathbf{a}_{t}^{T},b_{i})}{|I_{jk}|}+l_{jk}=|I_{jk}|(c_{ji}(\mathbf{a}_{i}^{T},b_{i})+\epsilon_{ji})+l_{jk}$
where $\sum_{j=1}^{m}l_{jk}^{2}=l_{k}^{2}$ the optimization criterion loss, and $\epsilon_{ji}$ is the error between $c_{ji}(\mathbf{a}_{i}^{T},b_{i})$ and the mean $\frac{\sum_{i\in I_{jk}}c_{ji}(\mathbf{a}_{t}^{T},b_{i})}{|I_{jk}|}$. It is easy to verify that $\epsilon_{ji}$ tends to zero as $\delta_{max}$ tends to 0.
@@
$\leq\sum_{k\in C_{j+}}\sum_{i\in I_{jk}^{\prime}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
@@
$\max_{\mathbf{u}\in V_{P_{j}}}\operatorname{dist}(u,\tilde{P}_{j})\leq\sum_{k\in C_{j+}}\sum_{i\in I_{jk}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
@@
$H(P_{j},\tilde{P}_{j})\leq\sum_{k\in C_{j+}}\sum_{i\in I_{jk}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
@@
$H(Q_{j},\tilde{Q}_{j})\\ \leq\sum_{k\in C_{j-}}\sum_{i\in I_{jk}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}\\ +\sum_{i\in N_{j-}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
@@
$\frac{1}{\rho}\max_{\mathbf{x}\in B}|v_{j}(\mathbf{x})-\tilde{v}_{j}(\mathbf{x})|\leq\sum_{k=1}^{m}\sum_{i\in I_{jk}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}\\ +\sum_{i\in N_{j}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
Using the fact that $I_{jk}\subset I_{k}$ and $N_{min}\leq|I_{jk}|,\forall j,k$ we have
$\frac{1}{\rho}\max_{\mathbf{x}\in B}|v_{j}(\mathbf{x})-\tilde{v}_{j}(\mathbf{x})|\leq\sum_{k=1}^{m}\sum_{i\in I_{k}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{N_{min}}+\|\epsilon_{ji}\|\right\}+\sum_{i\in N_{j}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
@@
$\sum_{j=1}^{m}|u_{j}|\leq\sqrt{m}\sqrt{\sum_{j=1}^{m}|u_{j}|^{2}}=\sqrt{m}\|(u_{1},\dots,u_{m})\|$
@@
$\sum_{j=1}^{m}|c_{ji}|\leq\sqrt{m}\|\mathbf{C}_{:,i}\|$
$\sum_{j=1}^{m}\|l_{jk}\|\leq\sqrt{m}\sqrt{\sum_{j=1}^{m}\|l_{jk}\|^{2}}=\sqrt{m}\cdot l_{k}$
$\sum_{j=1}^{m}\|\epsilon_{ji}\|\leq\sqrt{m}\sqrt{\sum_{j=1}^{m}\|\epsilon_{ji}\|^{2}}=\sqrt{m}\|\epsilon_{:,i}\|_{F}$

### v2

@@ A.6 Proof of Proposition 14
Consider any vertex $\mathbf{u}$ of $P_{j}$. This vertex can be written as the sum of generators $c_{ji}(\mathbf{a}_{i}^{T},b_{i})$, for some subset $I_{j+}^{\prime}\subseteq I_{j+}$. Thus,
$\mathbf{u}=\sum_{i\in I_{j+}^{\prime}}c_{ji}(\mathbf{a}_{i}^{T},b_{i}).$
For every generator $c_{ji}(\mathbf{a}_{i}^{T},b_{i}),i\in I_{j+}$ that belongs to positive cluster $k\in C_{j+}$, define
$x_{ji}=\arg\min_{x}\left\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})-x\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|,$
@@
$\sum_{j=1}^{m}\left\||\tilde{C}_{jk}|(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})-\sum_{i\in I_{jk}}|C_{ji}|(\mathbf{a}_{i}^{T},b_{i})\right\|^{2}.$
@@
By hypothesis, for every cluster $k$, the vectors of the set $\{(\mathbf{a}_{i}^{T},b_{i})|i\in I_{k}\}$ form pair-wise acute angles. Every vector of the set of vectors $\{\sum_{i\in I_{jk}}|c_{ji}|(\mathbf{a}_{i}^{T},b_{i})|j\in[m]\}$ lies inside the cone of the set of vectors $\{(\mathbf{a}_{i}^{T},b_{i})|i\in I_{k}\}$. It is easy to verify that the representative $|\tilde{c}_{jk}|(\tilde{\mathbf{a}}_{k},\tilde{b}_{k})$ that the iterative algorithm produces lies inside the cone of the set of vectors $\{\sum_{i\in I_{jk}}|c_{ji}|(\mathbf{a}_{i}^{T},b_{i})|j\in[m]\}$, which is a subset of the cone of the set of vectors $\{(\mathbf{a}_{i}^{T},b_{i})|i\in I_{k}\}$. Thus, the representative forms an acute angle with every vector of the set $\{(\mathbf{a}_{i}^{T},b_{i})|i\in I_{k}\}$, and thus $x_{ji}\geq 0$.
For every cluster $k\in C_{j+}$ define $I_{jk}^{\prime}=I_{k}\cap I_{j+}^{\prime}$ and
$\tilde{x}_{jk}=\sum_{i\in I_{jk}^{\prime}}x_{ji}.$
@@
Thus, for every cluster $k\in C_{j+}$, the point $\tilde{x}_{jk}\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ lies inside the segment $[\mathbf{0},\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$], and thus belongs to $\tilde{P_{j}}$.
For the vertex $\mathbf{u}$, we choose to compare it with the point
$\sum_{k\in C_{j+}}\tilde{x}_{jk}\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\in\tilde{P_{j}}.$
@@
$\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})=\sum_{i\in I_{jk}}c_{ji}(\mathbf{a}_{i}^{T},b_{i})+l_{jk}=|I_{jk}|\frac{\sum_{i\in I_{jk}}c_{ji}(\mathbf{a}_{t}^{T},b_{i})}{|I_{jk}|}+l_{jk}=|I_{jk}|(c_{ji}(\mathbf{a}_{i}^{T},b_{i})+\epsilon_{ji})+l_{jk},$
where $\sum_{j=1}^{m}\|l_{jk}\|^{2}=l_{k}^{2}$ the optimization criterion loss, and $\epsilon_{ji}$ is the error between $c_{ji}(\mathbf{a}_{i}^{T},b_{i})$ and the mean $\frac{\sum_{i\in I_{jk}}c_{ji}(\mathbf{a}_{t}^{T},b_{i})}{|I_{jk}|}$. It is easy to verify that $\epsilon_{ji}$ tends to zero as $\delta_{max}$ tends to 0.
@@
$\leq\sum_{k\in C_{j+}}\sum_{i\in I_{jk}^{\prime}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|.$
@@
$\max_{\mathbf{u}\in V_{P_{j}}}\operatorname{dist}(\mathbf{u},\tilde{P}_{j})\leq\sum_{k\in C_{j+}}\sum_{i\in I_{jk}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|.$
@@
$H(P_{j},\tilde{P}_{j})\leq\sum_{k\in C_{j+}}\sum_{i\in I_{jk}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}+\sum_{i\in N_{j+}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|.$
@@
$H(Q_{j},\tilde{Q}_{j})\\ \leq\sum_{k\in C_{j-}}\sum_{i\in I_{jk}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}\\ +\sum_{i\in N_{j-}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|.$
@@
$\frac{1}{\rho}\max_{\mathbf{x}\in B}|v_{j}(\mathbf{x})-\tilde{v}_{j}(\mathbf{x})|\leq\sum_{k=1}^{m}\sum_{i\in I_{jk}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{|I_{jk}|}+\|\epsilon_{ji}\|\right\}\\ +\sum_{i\in N_{j}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|.$
Using the fact that $I_{jk}\subseteq I_{k}$ and $N_{min}\leq|I_{jk}|,\forall j,k$ we have
$\frac{1}{\rho}\max_{\mathbf{x}\in B}|v_{j}(\mathbf{x})-\tilde{v}_{j}(\mathbf{x})|\leq\sum_{k=1}^{m}\sum_{i\in I_{k}}\min\left\{\|c_{ji}(\mathbf{a}_{i}^{T},b_{i})\|,\frac{\|l_{jk}\|}{N_{min}}+\|\epsilon_{ji}\|\right\}+\sum_{i\in N_{j}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|.$
@@
$\sum_{j=1}^{m}|u_{j}|\leq\sqrt{m}\sqrt{\sum_{j=1}^{m}|u_{j}|^{2}}=\sqrt{m}\|(u_{1},\dots,u_{m})\|.$
@@
$\sum_{j=1}^{m}|c_{ji}|\leq\sqrt{m}\|\mathbf{C}_{:,i}\|,$
$\sum_{j=1}^{m}\|l_{jk}\|\leq\sqrt{m}\sqrt{\sum_{j=1}^{m}\|l_{jk}\|^{2}}=\sqrt{m}\cdot l_{k},$
$\sum_{j=1}^{m}\|\epsilon_{ji}\|\leq\sqrt{m}\sqrt{\sum_{j=1}^{m}\|\epsilon_{ji}\|^{2}}=\sqrt{m}\|\epsilon_{:,i}\|_{F}.$